Holwood News

21/02/2010  TW Meeting
  HERA met with Taylor Wimpey on 16/02/2010 to discuss the outstanding works and snagging at Holwood. Residents were represented by Steve Turner (Kaz), Colin Glover, and Bill Markham. TW were represented by Beryl Foote (quantity surveyor)and Roy Taylor (project manager). The TW representatives are ex-employees of TW who were retained in November 2009 by TW to complete the outstanding works and snags here at Holwood. HERA first explained residents frustration at having met with numerous TW representatives in the past only to then see these representatives either be removed from office or be refused the funds by TW management to complete the agreed work. Whilst acknowledging that this was not the fault of BF and RT, HERA also complained at the apparent lack of progress since November and questioned the lack of supervision and quality control of operatives recently seen on site. BF and RT explained as follows:- They have been instructed by TW management to 'make residents happy with the development'.They are in possession of and are working to the 'outstanding works list' as produced by HERA. They are intending to deal with issues wherever possible to HERA satisfaction.They did not dispute any of the items HERA had identified as needing investigation/ attention. They will deal with each issue in turn and expect the works to take some months to complete. They expect to still be here at summertime. The delay so far was because of 'the weather' and by attempts to get 'original contractors' back on site to do remedial works under contractual obligation where possible. There had also been quite a good deal of investigative work. At the time of writing reported progress is as follows:- 1)water systems are being thoroughly cleaned and specialist contractors are instructed to commission the pumps in Block C. 2)a manual override for the water system is agreed so that we will have mains pressure in the event of electrical failure. 3) All rubbish/furniture has been collected from the various dumping places around the site, including the cycle store at Block C and is being taken away. There will be a delay whilst BF and RT now investigate the exact specification for the cycle store room. 4)the problem of 'flooding' on the podiums between the Blocks has been investigated and a solution is proposed. The corrective works are instructed and will begin in the next few weeks. 5)The roofs have been inspected and a number of defects identified. Works have started on corrective works but have been suspended during the recent wet weather.They will continue. 6) Much investigative work is still in progress and in the coming week one of the ground floor patios will be taken up to investigate the cause of water flooding over balconies instead of exiting via the drainage system as intended.This pilot is intended to reveal the solution for all balconies at Holwood. 7) BF has investigated the maingates, has condemned the electrical system, and has declared the rams inadequate for purpose.She has attempted to get the electrics replaced by the original installer (interphone) but this has proved unsuccessful. She has now asked 'Eagle Automation' to quote for complete new electrical installation and upgraded rams. This matter is now being treated as a matter of urgency. BF agreed to the incorporation of a timeclock in the new installation. This would allow HERA the option to consider keeping the gates open at peak traffic times. Further talks confirmed that BF was able to instruct works without further reference to anybody at TW but that she was reporting against a 'budget'. She was not willing to reveal the size of this budget. This is just a brief resume of lengthy talks. We agreed to meet and discuss progress on a monthly basis. The next meeting will be 23 March. HERA left the meeting encouraged that we are at last seeing progress, pleased at the openness of BF and RT, but with an ongoing concern that the quality of work and the size of budget will prove sufficient! This meeting will be further discussed in more detail by all Committe members on Tuesday 23 February. Kaz
  Article added by:Steve
  Go Back